### v2

# Security Code Review by Large Language Models

Abstract.
Security code review, as a time-consuming and labour-intensive process, typically requires integration with automated security defect detection tools to ensure code security. Despite the emergence of numerous security analysis tools, those tools face challenges in terms of their poor generalization, high false positive rates, and coarse detection granularity. A recent development with Large Language Models (LLMs) has made them a promising candidate to support security code review. To this end, we conducted the first empirical study to understand the capabilities of LLMs in security code review, delving into the performance, quality problems, and influential factors of LLMs to detect security defects in code reviews. Specifically, we compared the performance of 6 LLMs under five different prompts with the state-of-the-art static analysis tools to detect and analyze security defects. For the best-performing LLM, we conducted a linguistic analysis to explore quality problems in its responses, as well as a regression analysis to investigate the factors influencing its performance. The results are that: (1) existing pre-trained LLMs have limited capability in detecting security defects during code review but significantly outperform the state-of-the-art static analysis tools. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) GPT-4 makes few factual errors but frequently generates unnecessary content or responses that are not compliant with the task requirements given in the prompts. (4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic and written by developers with less involvement in the project.
Large Language Model, Code Review, Security Analysis
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/xxxxxxx.xxxxxxxisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Security defects can have severe consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004). Security code review is proven to be an effective process in the early stage of software development to prevent the introduction of security defects into code repositories, thereby mitigating the subsequent fixing and maintenance costs. Consequently, many popular projects and organizations have integrated security code review into their development pipelines (Paul et al., 2021). However, the resource-intensive nature of security code review, requiring significant human effort and time to review and revise code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024). Therefore, it would be helpful to develop effective automated tools to assist with detecting security defects during code review. Despite various security code analysis tools being proposed, they each encounter practical challenges such as high false positive rates, input range limitations, and weak scalability. Although data-driven approaches for security defect detection, especially those based on deep learning (DL), can somewhat handle these challenges, they inevitably lack robustness and tend to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024).
Pre-trained LLMs have demonstrated promising capabilities across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test improvement (Kang et al., 2023), and specification generation (Xie et al., 2023). Consequently, they are also considered a potential candidate for security code review. However, LLMs have not yet been employed in this specific field. Although many studies have examined the use of LLMs to detect security defects, most have been limited to quantitative evaluations of their performance (Bakhshandeh et al., 2023; Zhang et al., 2023; Zhou et al., 2024b), lacking in-depth analysis of specific problems encountered during the detection process and factors affecting the performance of LLMs in this task. These insights are crucial for enhancing the effectiveness of LLMs in security code review.
To bridge the knowledge gap, this study comprehensively explored the potential of LLMs in security code review from both qualitative and quantitative perspectives, including their capabilities, existing quality problems and influential factors. Specifically, We utilized a dataset of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator). First, based on the prompting strategies formulated by Zhang et al. (Zhang et al., 2023), we identified 5 prompt templates to enhance LLMs. Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers. After selecting the responses generated by the best-performing LLM-prompt combination, we first manually extracted quality problems from these responses. We then used these responses to construct a cumulative link mixed model to explore the impact of 10 factors on the performance of the LLM. Our findings reveal that: (1) current popular general-purpose LLMs exhibit limited capability in detecting security defects identified by reviewers in code reviews. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review. (4) GPT-4 is more likely to identify security defects in code files that have fewer tokens, contain functional logic and were written by developers with less involvement in the project. In summary, this study makes the following contributions:
(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review; (2) We firstly measured the impact of the randomness of LLMs on the consistency of security code review; (3) We identified quality problems in the responses generated by the top-performing LLM to present the challenges it faces in security code review; (4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
2. Related Work
Several studies have explored the performance of LLMs in detecting security defects. The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type. Zhou et al. utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgement under various prompts (Zhou et al., 2024b). They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art security defect detection approaches, while GPT-4 consistently outperformed the state-of-the-art. Purba et al. compared the performance of LLMs with popular static analysis tools in binary detection. Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools (Purba et al., 2023). Unlike previous studies, our work provides a more granular evaluation of the capability of LLMs in file-level security code reviews, by including additional insights such as defect localization, description, and possible solutions.
Previous studies also investigated whether LLMs could provide more information in security defect detection. Bakhshandeh et al. provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers(Bakhshandeh et al., 2023). Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools. Yin et al. evaluated LLMs’ performance across multiple tasks related to security defects, including defect detection, localization, and description (Yin and Ni, 2024). They found that the existing state-of-the-art approaches such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023) generally outperformed LLMs. Specifically, in the tasks of defect localization and description, LLMs exhibited limited overall accuracy and varied performance across different CWE types. Most of the aforementioned studies utilized datasets that were synthesized or collected from tool documentation and CVE databases, while our study adopted a dataset derived from real-world code reviews in four industrial open-source projects, thus providing a reflection of the practicability of LLMs in security code review.
Employing specific prompting frameworks to enhance the performances of LLMs in identifying security defects has also been explored. Yang et al. introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection (Yang et al., 2024). Their framework yielded superior performance compared to state-of-the-art prompting techniques. Nong et al. proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information, to evaluate the performance of LLMs in identifying security defects (including their classification and resolution) (Nong et al., 2024). The main differences between the two studies above and our study are that: (1) Yang et al. and Nong et al. employed CoT through few-shot learning, whereas we used multi-turn dialogue for CoT. (2) Nong et al. utilized synthetic datasets and CVE datasets, whereas we utilized a manually curated security code review dataset obtained from open-source projects to guarantee data quality and reflect real-world applications. (3) Nong et al.’s work covered five CWE categories, whereas the dataset in our work included 15 security defect types, which are considered the most frequently identified types in code reviews (Yu et al., 2023).
In contrast to all the aforementioned studies, besides quantitatively evaluating LLM performance in security defect detection, we conducted more in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance. Thus, we are able to gain a more comprehensive understanding of the potential and challenges faced by LLMs in security code review.
3. Methodology
3.1. Research Questions
The goal of this work is to comprehensively assess the capability of LLMs in assisting with security code review, by automatically detecting security defects in the given code. To achieve this goal, we divided our study into three Research Questions (RQs), as shown below. The research procedure of each RQ is illustrated in Fig. 1.
RQ1: How do LLMs perform in detecting security defects identified by reviewers in code reviews?
RQ1.1: How do LLMs perform in security code review under the basic prompt?
RQ1.2: Can prompting with auxiliary information improve the performance of LLMs?
RQ1.3: Can the CoT prompting approach improve the performance of LLMs?
RQ1.4: What is the impact of the randomness of LLMs on the consistency of security code review?
Motivation: Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a). To address these limitations, RQ1 utilizes a manually-curated dataset to ensure data quality and comprehensively evaluates the capability of LLMs in reviewing a complete code file from real-world code reviews for security defects. Additionally, due to the inherent randomness of LLMs, for the same prompt, the responses of LLMs may exhibit considerable divergence. However, this aspect of LLMs has not yet been studied when applied to security tasks. Hence, we conducted a quantitative analysis of the consistency of LLM responses to explore the impact of LLM randomness on the task of security code review.
RQ2: What are the quality problems in LLM-generated responses when applied to detect security defects?
Motivation: Kabir et al. noted that the responses of ChatGPT to programming questions may be verbose, inconsistent, and contain conceptual or logical errors (Kabir et al., 2024), which negatively impact the quality of the answers. However, for security code review, the quality problems present in responses generated by LLMs have not been thoroughly analyzed in previous studies. Hence, RQ2 conducts an in-depth linguistic analysis of LLM responses for quality problems, and examines the types and distribution of these problems to bridge the knowledge gap and indicate potential areas for improvement.
RQ3: Which factors influence the performance of LLMs in security code review?
Motivation: Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (Chen et al., 2023; Kabir et al., 2024; Zhou et al., 2023), primarily through case studies or observations. However, there is a lack of evidence substantiating such claims. RQ3 employs a regression analysis to validate whether these factors have a significant impact on the performance of LLMs. This RQ helps to pinpoint areas of concerns in applying LLMs to security code review and highlight possible directions for improvement.
Figure 1. An overview of the research procedure for investigating the three RQs
3.2. Data Collection
We leveraged the code review dataset constructed by Yu et al. (Yu et al., 2023) and examined the capability of LLMs in security code review. The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types, from four projects: OpenStack Nova and Neutron, and Qt Base and Creator. These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018). For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs. Instead, we included only the code from the single file, where a review comment was made by the reviewer to identify a security defect. The 614 security-related review comments were reorganized as follows:
First, we filtered out comments that were published at the patchset level, as such data cannot be located in a specific file to construct a prompt. Then, for each code file that these comments correspond to, we consolidated all security-related comments in the file and the corresponding security defect records into a single data item. Next, we utilized Gerrit API to retrieve the content of each code file and the commit messages of the patchsets they belong to. During this procedure, if we cannot obtain the content of the code files through API, we ignored those files. Lastly, since 97% files in this dataset are Python and C/C++ files, we excluded those written in languages other than these two (e.g., .json and .sh).
The final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
3.3. Research Procedure of RQ1
3.3.1. LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December, 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through API. For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs. Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings. The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform. During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways. The two models provided by the ChatGPT platform performed better, thus were employed in our study.
Table 1. Versions and hyper-parameters of 6 LLMs studied
Model Version Context Length Parameter
temperature top_p
GPT-3.5 May 2024 version - -
GPT-4 May 2024 version - -
GPT-4 Turbo gpt-4-1106-preview 128k 1.0 1.0
Gemini Pro gemini-1.0-pro 32k 0.9 1.0
Llama 2 7B llama-2-7b-chat 4k 1.0 0.9
Llama 2 70B llama-2-70b-chat 4k 1.0 0.9
3.3.2. Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (Zhang et al., 2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected five best-performing prompts. The detailed construction strategies of the five prompts are provided in (Anonymous, 2024).
  1. (1)
    Prompt 1 ($\textbf{P}_{b}$): the basic prompt. We directly instruct the LLM to review the provided code for security defects and output their description, code line numbers, and fixing solution. If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
  2. (2)
    Prompt 2 ($\textbf{P}_{r}$): $\textbf{P}_{b}$ + Project Information. $\textbf{P}_{r}$ is designed based on $\textbf{P}_{b}$ but includes the name of the source project as auxiliary information. The LLM is asked to act as a code reviewer for the project of the provided code in the prompt.
  3. (3)
    Prompt 3 ($\textbf{P}_{c}$): $\textbf{P}_{b}$ + General CWE Instruction. $\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects, without providing specific CWE details.
  4. (4)
    Prompt 4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction. $\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (Pan et al., 2023)). It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries. The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$)..
  5. (5)
    Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$. The task of security code review is split into two steps. In $\textbf{P}_{cot-1}$, given that we only provide a single code file from a complete project in the prompt, we include the corresponding commit message of the file and instruct the LLM to generate code context according to the commit message. Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
3.3.3. Evaluation Metrics
Aiming to measure the performance of LLMs, we adopted the evaluation methodology proposed by Mahajan et al. (Mahajan et al., 2020) to rate the responses that LLMs generate with some adjustments, as detailed in Table 2. First, we excluded the ‘Unavailable’ category because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers, while a lower M-Score signifies less misleading results produced during security defect detection. Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness, and calculated the average of the three performance scores as the final result.
Table 2. Adjusted IHMU-category for LLM responses
Classification Definition
Instrumental (I) The response explicitly indicates the existence of the security defect identified by the reviewer and provides a fully accurate description.
Helpful
The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
Uncertain
The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
3.3.4. Baselines
We attempted different tools and learning techniques for detecting security defects to select appropriate baselines for our work. To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Cppcheck Solutions AB, 2024) for C/C++ and Bandit (ban, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SonarSource SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like python, but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (sem, 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. In order to obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
3.3.5. Data Labelling
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. To rate the responses, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then the first and third authors proceeded to label all the rest responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset and found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 4.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
3.3.6. Consistency Calculation
To measure the randomness of LLMs, we adopted the metrics in Chang et al.’s work (Chang et al., 2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, for a given code file under a specific prompt template, an LLM generated three responses. We then calculated the entropy for the four response categories in the three responses and determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
3.4. Research Procedure of RQ2
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze quality problems present in these responses. We employed the open coding and constant comparative method (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
Based on the problem categories formulated by Kabir et al. (Kabir et al., 2024) and problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. Notably, we excluded the Consistency theme defined in Kabir et al. (Kabir et al., 2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, 10 responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, problem analysis is a multi-level and multi-label classification task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the remaining data analysis was completed by the two authors and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each problem type, we computed its average frequency in three repetitive experiments as the final occurrence rate of such problem in GPT-4 responses.
3.5. Research Procedure of RQ3
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the randomness of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is outlined in the following subsections.
3.5.1. Attributes Construction
To comprehensively collect factors influencing the detection of security defects by LLMs, excluding the prompt and the model itself, we undertook the following steps:
First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work. Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntoshet al.’s work (McIntosh et al., 2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens. Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al.’s work  (Chen et al., 2023), transforming them into factors for our experiment. Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations. In the end, we obtained a list of 10 factors in Table 3, along with the source, definition, and rationale for each factor.
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments and obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment, along with their factor values, ready for model fitting.
Table 3. Factors that may influence the performance of LLM in security code review
Factor Source Definition Rationale
Token
Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file. Computed using the TikToken (tik, 2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
FileType
Created Type of the code file, i.e., Source or Auxiliary. Source indicates files that directly contain the logic of the application (i.e., .cpp, .c and .py) while Auxiliary files are utilized for declarations (i.e., .h and .hpp). The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
SecurityDefectType
Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. (Yu et al., 2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit
Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity
Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976). Computed using the Lizard (liz, 2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
Community
Created Name of the community source of the code file, i.e., OpenStack or Qt. OpenStack and Qt differ significantly in terms of their structure, technology stack, project function, and main components. The LLMs’ performance may differ on data derived from these two communities.
AuthorExperience
Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
AnnotationRatio
Created The ratio of manual annotations in the code file. If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
AnnotationHasCode
Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.
AnnotationisSecurityRelated
Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects). Keywords in the description related to target security defect in annotations can act as cues for LLMs, while the descriptions of non-target security defects often involve the historical fixing or security mechanisms, which may cause interference to LLMs.
3.5.2. Regression Model Construction
Correlation and Redundancy Analysis: If explanatory variables are highly correlated with each other, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 2. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 2, where each factor is denoted by an abbreviation, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, which can distort the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Figure 2. The Phi_K correlation coefficients of explanatory variables.
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman rank correlations (${\rho}^{2}$) between them. As recommended by Harrell  (Harrell, 2001), explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5. According to the computation results shown in Fig. 3, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘TK’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. For categorical variables, they should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
Figure 3. Dotplot of the Spearman ${\rho}^{2}$ of explanatory variable and response variable in three repetitive experiments
Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015). To balance model complexity and fit, we tested models with and without allocating 3 degrees of freedom to the ‘TK’ factor. The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Anonymous, 2024).
4. Results and Analysis
4.1. RQ1 Performance of LLMs
Based on the evaluation metrics defined in Section 3.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts and compared the performance of LLMs with baseline tools in the C/C++ and Python dataset, respectively. If two out of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the other ones. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: 1) the response count and performance scores of each LLM-prompt combination on the complete dataset (Table 4), and 2) the performance scores of each LLM-prompt combination on the selected 57 instances (Table 5).
From Table 4, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 5 demonstrated that in the selected 57 cases, Llama 2 70B scored obviously lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in Table 4. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code review are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools.
Table 4. Performance scores and entropy for each LLM-Prompt combination on the python and c/c++ dataset
Python C/C++
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 Resp. Count 123 123 123 123 122 GPT-3.5 Resp. Count 141 141 141 139 137
I-Score 1.08% 1.35% 0.27% 0.27% 1.37% I-Score 1.89% 2.13% 0.00% 0.24% 0.73%
IH-Score 2.71% 3.80% 1.35% 0.27% 5.19% IH-Score 5.67% 3.31% 0.47% 0.96% 1.46%
M-Score 89.37% 84.28% 92.41% 97.29% 69.12% M-Score 76.60% 86.05% 99.29% 95.44% 84.43%
Entropy 0.2572 0.3692 0.2070 0.0746 0.6842 Entropy 0.4860 0.2439 0.0195* 0.1153 0.2931
GPT-4 Resp. Count 175 175 175 174 174 GPT-4 Resp. Count 247 247 247 247 247
I-Score 4.76% 3.43% 3.24% 5.17*3.06% I-Score 4.45% 3.91% 3.91% 5.40% 5.26%
IH-Score 9.34% 9.52% 8.38% 14.56*8.62% IH-Score 7.96% 8.01% 9.31% 11.74% 11.47%
M-Score 61.31% 60.95% 46.47% 44.06% 61.69% M-Score 55.47% 64.24% 44.53% 39.00% 43.59%
Entropy 0.7426 0.7283 0.8600 0.9416 0.7450 Entropy 0.6215 0.5329 0.6505 0.6955 0.7330
GPT-4 Turbo Resp. Count 201 201 201 201 201 GPT-4 Turbo Resp. Count 269 269 269 269 269
I-Score 0.66% 0.33% 0.66 0.83 0.17 I-Score 0.25% 0.12% 0.25% 0.50% 0.25%
IH-Score 0.66 0.83 2.32 1.99 0.33 IH-Score 0.25% 0.37% 1.24% 1.61% 1.12%
M-Score 97.18% 96.35% 85.74% 92.87% 98.01% M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
Entropy 0.1757 0.0746 0.3370 0.6125 0.0404* Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
Gemini Pro Resp. Count 162 161 160 160 159 Gemini Pro Resp. Count 223 221 221 221 220
I-Score 0.00% 0.00% 0.63% 1.25% 0.21% I-Score 0.00% 0.15% 0.30% 1.66% 0.00%
IH-Score 0.41% 0.00% 0.84% 3.54% 0.21% IH-Score 0.45% 0.15% 0.90% 3.77% 0.45%
M-Score 90.54% 97.09% 84.04% 55.21% 98.53% M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
Entropy 0.0777 0.0901 0.2424 0.1873 0.0548 Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
Llama 2 7B Resp. Count 62 62 61 61 50 Llama 2 7B Resp. Count 57 57 57 51 33
I-Score 0.00% 0.54% 1.62% 2.19% 1.33 I-Score 0.58% 0.00% 0.00% 2.61% 1.01%
IH-Score 1.61% 1.61% 5.40% 8.74% 4.67% IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
M-Score 87.10% 83.33% 56.78% 40.98% 33.33*M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
Entropy 0.2625 0.3070 0.4766 0.4351 0.5521 Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
Llama 2 70B Resp. Count 62 62 62 61 47 Llama 2 70B Resp. Count 57 57 57 51 37
I-Score 0.54% 1.07% 2.15% 3.83% 1.42% I-Score 1.17% 1.17% 1.75% 6.54*1.80%
IH-Score 4.84% 4.30% 6.45% 11.48% 3.55% IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
M-Score 79.57% 84.41% 58.60% 43.17% 36.17% M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
Entropy 0.4470 0.3474 0.3514 0.3832 0.5762 Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
Baselines SAST SonarQube CodeQL Bandit Semgrep Baselines SAST CppCheck Semgrep
Resp. Count 258 258 258 258 Resp. Count 276 276
I-Score 0.00% 0.39% 0.00% 0.00% I-Score 0.00% 0.00%
IH-Score 0.00% 1.16% 1.16% 0.00% IH-Score 0.72% 0.36%
M-Score 91.47% 82.56% 78.29% 90.31% M-Score 98.19% 96.74%
  • The best values for each score are highlighted in bold and marked with *.
Table 5. Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 I-Score 3.51% 3.51% 0.58% 0.58% 2.34%
IH-Score 7.60% 8.77% 1.17% 0.58% 4.09%
M-Score 81.87% 79.53% 94.74% 95.32% 77.19%
GPT-4 I-Score 8.77% 8.77% 5.85% 12.87*9.94
IH-Score 20.46% 19.30% 14.04% 23.98*19.88
M-Score 43.86% 42.11% 45.61% 40.35% 41.52%
GPT-4 Turbo I-Score 0.58% 1.75% 2.34% 2.34% 0.00%
IH-Score 0.58% 3.51% 5.26% 4.09% 1.75%
M-Score 97.08% 91.23% 75.44% 88.30% 92.98%
Gemini Pro I-Score 0.58% 0.00% 0.00% 0.58% 0.00%
IH-Score 1.17% 0.00% 0.58% 2.92% 0.00%
M-Score 96.49% 100.00% 88.30% 76.61% 100.00%
Llama 2 7B I-Score 0.58% 1.17% 1.17% 0.58% 1.75%
IH-Score 1.75% 2.34% 4.09% 2.34% 7.60%
M-Score 87.13% 84.21% 60.82% 53.80% 25.73%
Llama 2 70B I-Score 2.34% 1.17% 2.92% 5.26% 4.09%
IH-Score 4.68% 2.34% 7.02% 17.54% 4.68%
M-Score 70.18% 80.12% 42.69% 20.47*44.44%
4.1.1. RQ1.1
When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%). Following closely behind is Llama 2 70B and GPT-3.5. GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit. Similar results are observed in Table 5. Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, GPT-4 achieved I-Score = 8.77%, IH-Score = 20.46%, and M-Score = 43.86%. It is evident that GPT-4’s performance on the selected 57 instances exceeds its performance across the entire dataset. Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
4.1.2. RQ1.2
As shown in Table 4, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small. Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively. Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement. A possible reason is that the name of components or packages utilized may already imply the source project of the code file, thereby diminishing the impact of project information provided in $\textbf{P}_{r}$. Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific. We speculate that some LLMs may already understand the task described in the prompt quite well, thus introducing a persona could add unnecessary complexity of the prompt and lead to distractions.
We can see from Table 4 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects, but it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, how to leverage CWE information in the prompt is a question worthy of consideration.
4.1.3. RQ1.3
Table 4 illustrates that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B. However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs. As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets. Such instability may be attributed to differences in training data among different LLMs. Additionally, we first instructed the LLM to generate missing code context based on the corresponding commit message, where the consistency between this generated context and the actual code context cannot be guaranteed. Through analyzing specific cases, we discovered that some LLMs automatically fixed security defects while generating code context, leading to confusing and erroneous detection results in the second step of detecting security defects in the provided code.
4.1.4. RQ1.4
According to the mean entropies of various LLM-prompt combinations in Table 4, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given the poor security defect detection capabilities of these two LLM-prompt combinations, their high consistency may be due to their tendency to not detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review, but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with stronger security defect detection capabilities may also exhibit greater randomness.
4.2. RQ2 Quality Problems in Responses
The average distribution of 12 problem types across 4 themes in responses to three repetitive experiments is illustrated in Fig. 4, where the same color represents types within the same theme, with darker shades indicating the themes.
Figure 4. Average proportion of each problem type present in responses generated by GPT-4
From the perspective of the four themes, a vast majority of responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
Correctness: The responses reveal two types of incorrectness: Incorrect fact (6.25%) and Incorrect concept (10.13%). The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario. Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type. As shown in the example below, the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
Example 1 - Incorrect concept
GPT-4: ...**CWE-682 (Incorrect Calculation):**
**Issue:** The method _inventory_has_changed uses a comparison logic that might not be entirely accurate in all scenarios...
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to the situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing overly general solutions. An example of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
Example 2 - Vague statement
GPT-4: ...
**CWE-703(Improper Check or Handling of Exceptional Conditions)**
**General Concern**: Throughout the code, there are instances where error conditions might not be handled comprehensively...
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appear in the responses generated by GPT-4. Such information includes the reasons for the absence of security defect in the current code, as well as general suggestions on security measures and review processes. As illustrated in the example below, the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence is deemed to be Unnecessary information.
Example 3 - Unnecessary
GPT-4: ...**CWE-284 (Improper Access Control)**:
- There doesn’t seem to be any direct issue related to improper access control in the given code...
Compliance: Three types of non-compliance were identified with regard to whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and solutions of detected security defects requested in the prompt. From the example below, we can see that while the potential memory management issues caused by temporary files and directories are described and a solution is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
Example 4 - Incomplete answer
GPT-4: ### 3. CWE-664
**Potential Issue:** The creation of temporary files and directories (e.g., in function createQmlrcFile‘) ... can potentially lead to resource management issues.
**Solution:** Implement thorough checks and error handling...
4.3. RQ3 Factors Influencing LLMs
4.3.1. Model Stability
To assess the adequacy of our regression model’s fit, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation present in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
4.3.2. Explanatory Power of Each Factor
The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable. A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model. We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact. As shown in Table 6, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (¡.0001). Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$¡.0001) and ($\chi^{2}$=18.28, $p$¡.0001). Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively. However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold. Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom. As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
Table 6. Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(¿Chi)
TK 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
CMT Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
FT Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
ASR -1:0 0.97 2 9.95 0.0069**
1:0 2.69
AE 0.87 1 4.89 0.0271*
CT 1.01 1 3.06 0.0801
SDT Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
Crash:Race Condition 0.86
Deadlock:Race Condition 0.92
DoS:Race Condition 0.87
Encryption:Race Condition 1.18
Improper Access:Race Condition 0.68
Integer Overflow:Race Condition 0.86
Resource Leak:Race Condition 1.41
Use After Free:Race Condition 0.37
Extra:Race Condition 1.43
CPT 1.03 1 1.74 0.1868
AR 0.56 1 1.55 0.2126
AHC 1.11 1 1.09 0.2973
  • Statistical significance of explanatory power according to Wald $\chi^{2}$ test:
    ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$
4.3.3. Relationship between Factors and Responses
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 6). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files that contain functional logic, have fewer tokens, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that the LLM may be disrupted by annotations and less likely to detect the target security defect.
5. Implications
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (Steenhoek et al., 2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing the accuracy of detection. However, in real-world security reviews, although with the experience and domain knowledge, the reviewer is not guaranteed to consider all potential security defect types in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al.’s work (Bakhshandeh et al., 2023), providing a list of all 75 CWEs in their prompts resulted in a decrease in the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to effectively focus on the target defect, thereby yielding adverse effects.
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, it appears that annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations there are in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend developers write code comments in a standardized way to clarify the purpose of each comment, thus better support LLMs in code review tasks.
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amount of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs to overlook certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo has a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and LLama 2. Therefore, how to ensure the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
6. Threats to Validity
Internal Validity: One threat could be data leakage. The dataset used in our study covers two large-scale open-source projects, which may overlap with the data used for training LLMs. Since the training data time of GPT-3.5 and GPT-4 is not clearly disclosed, we cannot avoid this overlap by filtering the data for a specific time period. Moreover, according to the findings by Cao et al. (Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cut-off date. Another threat is the design of prompts. To evoke the best model performance, we design our prompt according to the best practices. We selected words that are as unbiased, not suggestive, and unambiguous as possible and iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. Lastly, LLMs may generate varied responses for the same prompts due to their inherent randomness. To mitigate this threat, we conducted three repetitive experiments for RQ1. For RQ2, we calculated the average frequency of each problem across these three experiments. In RQ3, we combined the responses from the three experiments to fit the model.
External validity: In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale). In dataset collection, we used a dataset of code reviews collected from real-world open-source projects that are widely used in code review-related studies. We acknowledge that the dataset may not fully represent industrial projects, and projects from proprietary software development are needed to increase the generalizability of the results.
Construct validity: With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure. These models are also not made publically available via API, thereby hindering the reproducibility of this study. Therefore, we explicitly indicate the versions of these two models used in the experiments, and provided all the responses generated by these models (Anonymous, 2024). In addition, the data labeling and extraction processes in this study were manually conducted, which may introduce subjective judgements. Thus, we adopted an open coding and constant comparative approach, conducting pilot experiments before all formal manual tasks to ensure a consensus among multiple authors.
7. Conclusions
In our study, we conducted a comprehensive analysis of the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems present in responses generated by this LLM and performed regression analysis to explore factors influencing its performance. Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to the state-of-the-art static analysis tools. However, their application remains limited. (3) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) During code review, GPT-4 is more effective in identifying security defects in code that has fewer tokens, containing functional logic and written by developers with less involvement in the project. (4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.