### v4

# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

Abstract
Security code review is a time-consuming and labor-intensive process typically requiring integration with automated security defect detection tools. However, existing security analysis tools struggle with poor generalization, high false positive rates, and coarse detection granularity. Large Language Models (LLMs) have been considered promising candidates for addressing those challenges. In this study, we conducted an empirical study to explore the potential of LLMs in detecting security defects during code review. Specifically, we evaluated the performance of six LLMs under five different prompts and compared them with state-of-the-art static analysis tools. We also performed linguistic and regression analyses for the best-performing LLM to identify quality problems in its responses and factors influencing its performance. Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review but significantly outperform the state-of-the-art static analysis tools. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) GPT-4 frequently generates verbose or non-compliant responses with the task requirements given in the prompts. (4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic, or written by developers with less involvement in the project.
keywords:
Large Language Model, Code Review, Security Analysis
1 Introduction
Security defects are potential security risks or weaknesses in the system introduced during implementation (Paul et al., 2021). Once attackers maliciously exploit a security defect, it could lead to serious consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004). The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (Planning, 2002). Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (GitLab, 2022). Code review is a human-intensive process in which the reviewer examines the code submitted by the developer to detect bugs, verify the implementation of the specification, ensure compliance with guidelines, and ensure quality. During this process, the reviewer raises issues in the code, discusses them with the developer, and provides recommendations. Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (Weir et al., 2017), from being merged into the source code repository. In this study, we focus on security analysis in the context of code review, referring to it as security code review (Edmundson et al., 2013).
Security code review is increasingly integrated into the development pipelines by project teams (Paul et al., 2021). Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024). Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review (Gholami and Amri, 2021). Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013), and weak scalability (Nong et al., 2021).
Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (Chakraborty et al., 2021). However, due to the lack of sizable and realistic training datasets (Nong et al., 2022), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024a). Nong et al. (2024b) found that even using the training dataset augmented by the latest vulnerability injection technique, the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively. Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test generation (Kang et al., 2023; Schäfer et al., 2023), and specification generation (Xie et al., 2023). Such models can understand and generate explanatory natural language responses, and are thus considered a potential candidate for security code review. We expect that LLM can act as a code reviewer, not only identifying security defects in code files, but also explaining the specific defect scenarios in detail to developers in a human-readable way. Applying LLMs to support security code review can help detect security defects earlier and improve productivity by helping code reviewers check their code faster and more efficiently. However, applying LLMs in the specific field of security code review remains largely unexplored.
Many studies have focused on leveraging LLMs for vulnerability detection. Zhou et al. (2024b) evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts, while Purba et al. (2023) compared LLMs with static analysis tools in detecting software vulnerabilities. These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks to judge whether there were security defects in the code. They did not require LLMs to provide more detailed information, such as the location and type of vulnerability. Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by Yang et al. (2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al. (2024a). However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
Motivated by the limitations mentioned above, we aim to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories. Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description, and suggested fix. The dataset we utilized consists of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator). We evaluated the performance of LLMs, analyzed existing quality problems, and examined factors influencing their performance to offer insights into the real-world applicability and limitations of LLM in security code review. First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by Zhang et al. (2023)). Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers. After selecting the responses generated by the best-performing LLM-prompt combination, we first manually extracted quality problems from these responses. We then used these responses to construct a cumulative link mixed model to explore the impact of 10 factors on the performance of the LLM. Our findings reveal that: (1) current popular general-purpose LLMs exhibit limited capability in detecting security defects identified by reviewers in code reviews. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review. (4) GPT-4 is more likely to identify security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project. In summary, this study makes the following contributions:
(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review. (2) We measured the impact of the non-determinism of LLMs on the consistency of security code review. (3) We identified quality problems and quantified hallucinations in the responses generated by the top-performing LLM, aiming to highlight the challenges faced by LLMs in security code review. (4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
Paper Organization. Section 6 surveys the related work. Section 2 describes the methodology employed in this study. Section 3 presents the results, followed by a discussion of the implications of our study results in Section 4. Section 5 clarifies the threats to validity. Finally, Section 7 concludes this work and outlines future directions. The replication package of this study has been provided online (Yu et al., 2025).
2 Methodology
2.1 Research Questions
This work aims to comprehensively assess the capability of LLMs to assist with security code review by automatically detecting security defects in the given code. We divided our study into three Research Questions (RQs), as shown below. The research procedure for each RQ is illustrated in Fig. 1.
RQ1: How effectively do LLMs detect security defects during code review?
RQ1.1: How effectively do LLMs perform security code reviews under the basic prompt?
RQ1.2: Can prompting with auxiliary information improve the performance of LLMs?
RQ1.3: Can the Chain-of-Thought (CoT) prompting approach improve the performance of LLMs?
RQ1.4: What is the impact of the non-determinism of LLMs on the consistency of security code review?
Motivation. Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a). To address these limitations, RQ1 utilizes a manually curated dataset to ensure data quality. It comprehensively evaluates the capability of LLMs to review a complete code file from real-world code reviews for security defects. Additionally, due to the inherent non-determinism of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt across different runs. Non-determinism is mainly caused by decoding configurations (Song et al., 2024), with other hypotheses such as floating point, unreliable GPU calculations (Ouyang et al., 2025), and its sparse MoE architecture failing to enforce per-sequence determinism (La Malfa et al., 2024; Puigcerver et al., 2023). Considering that the non-determinism of LLMs has not yet been studied when applied to security tasks, we conducted a quantitative analysis of the consistency of LLM responses to explore its impact on the task of security code review.
RQ2: What quality problems exist in LLM-generated responses during security code review?
RQ2.1: What is the distribution of quality problems in these responses?
RQ2.2: Which of these quality problems are hallucination-related, and what are their proportions?
Motivation. Kabir et al. (2024) noted that LLMs’ responses may be verbose, inconsistent, and contain conceptual or logical errors, negatively impacting the quality of the answers. Among these problems, hallucination in particular has been extensively studied  (Liu et al., 2024; Tonmoy et al., 2024). Hallucination refers to content generated by LLMs that appears plausible but lacks factual grounding (Huang et al., 2025). If we applied LLMs to real-world software development tasks, particularly during security code review, such hallucination may mislead practitioners, increase review costs, and even compromise system security. Previous studies have not yet thoroughly examined quality problems present in LLM-generated responses during security code review. RQ2 intends to conduct an in-depth linguistic analysis to investigate possible hallucination issues. First, we investigated the distribution of quality problems in LLMs’ responses, such as Vague statement, Incorrect concept, and Incomplete answer. Then we identified hallucination-related issues among these problems and quantified their proportion to bridge the knowledge gap and highlight potential areas for improvement.
RQ3: Which factors influence the performance of LLMs in security code review?
Motivation. Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (Chen et al., 2023; Kabir et al., 2024; Zhou et al., 2023), primarily through case studies or observations. However, there is a lack of evidence substantiating such claims. RQ3 employs a regression analysis to validate whether these factors significantly impact the performance of LLMs. This RQ helps pinpoint areas of concern in applying LLMs to security code review and highlights possible improvement directions.
Figure 1: An overview of the research procedure for investigating the three RQs
2.2 Data Collection
We leveraged the code review dataset constructed by Yu et al. (2023) and examined the capability of LLMs in security code review. The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (almost entirely written in Python), and Qt Base and Creator (mostly in C++ with some portions in C, Objective-C++, and Python). These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018). For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs. Instead, we included only the code from the single file, where the reviewer commented to identify a security defect. The 614 security-related review comments were reorganized as follows:
First, we filtered out comments published at the patchset level, as such data cannot be located in a specific file to construct a prompt. Then, for each code file to which these comments correspond, we consolidated all security-related comments in the file and the corresponding security defect records into a single data item. Next, we utilized the Gerrit API to retrieve the content of each code file and the commit messages of the patchsets to which they belong. During this procedure, if we cannot obtain the content of the code files through the API, we ignore those files. Lastly, since 97% of the files in this dataset are Python and C/C++ files, we decided to remove any other files (all are configuration and scripting files) as those only formed $\approx 3\%$ of the files and thus had a negligible impact on our results.
Our final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
2.3 Research Procedure of RQ1
Table 1: Versions and hyper-parameters of 6 LLMs studied.
Model Version Ctx Len Parameter
temperature top_p
GPT-3.5 May 2024 version - -
GPT-4 May 2024 version - -
GPT-4 Turbo gpt-4-1106-preview 128k 1.0 1.0
Gemini Pro gemini-1.0-pro 32k 0.9 1.0
Llama 2 7B llama-2-7b-chat 4k 1.0 0.9
Llama 2 70B llama-2-70b-chat 4k 1.0 0.9
2.3.1 LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API. For open-source LLMs, we deployed the smallest and largest chat versions of models in the Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs, each with 24 GB of VRAM. To realistically reflect on the practical capabilities of LLMs in security defect detection, we used the default hyperparameter settings throughout our experiments. Some studies mitigate the non-determinism of LLMs by setting the temperature to zero (Atil et al., 2024; Blackwell et al., 2024). In contrast, we refrained from doing so, as this may adversely affect the performance of the LLMs (Song et al., 2024) and may not guarantee determinism (Ouyang et al., 2025; Astekin et al., 2024). We retained the default temperature settings and quantified the non-determinism of LLMs as one aspect of performance evaluation in RQ1.4, rather than artificially suppressing such non-determinism via hyperparameter tuning. The versions, context lengths (Ctx Len), and parameter settings of LLMs used in our study are summarized in Table 1.
Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform. During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways. The two models provided by the ChatGPT platform performed better and thus were employed in our study.
2.3.2 Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected the five best-performing prompts. Fig. 2 illustrates the construction strategies for the five prompts.
  • Prompt 1 ($\textbf{P}_{b}$): the basic prompt. We instruct the LLM to review the provided code for security defects and output their description, code line numbers, and suggested fix. If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
  • Prompt 2 ($\textbf{P}_{r}$): $\textbf{P}_{b}$ + Project Information. $\textbf{P}_{r}$ is designed based on $\textbf{P}_{b}$ but includes the name of the source project as auxiliary information. The LLM is asked to act as a code reviewer for the project using the provided code in the prompt.
  • Prompt 3 ($\textbf{P}_{c}$): $\textbf{P}_{b}$ + General CWE Instruction. $\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects without providing specific CWE details.
  • Prompt 4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction. $\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (Pan et al., 2023)). It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries. The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$). We acknowledge that providing a list of CWEs may limit the scope of detection set of CWEs and miss security defects that do not fall into the provided CWE categories. However, including all CWEs is infeasible due to the limited context window of the LLM and the potential loss of focus.
  • Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$. Since we only included the code of a single file in the prompt instead of an entire patchset, we designed $\textbf{P}_{cot}$ to analyze the impact of missing contextual information on the performance of LLMs. The task of security code review is split into two steps. In $\textbf{P}_{cot-1}$, given that we only provide a single code file from a complete project in the prompt, we include the corresponding commit message of the file and instruct the LLM to generate code context according to the commit message. Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
Figure 2: Construction templates for the five prompts
Table 2: Adjusted IHMU-category for LLM responses.
Classification Definition Instance
Instrumental (I) The type of the security defect identified in the response is exactly the same as the target security defect, the location is strictly within the code line range of the target security defect, and the defect description can provide effective information for understanding the scenario. Reviewer: (Line 76)I think this is race-y. Let’s say we have a three coroutines scheduled to call function. First one gets to L74 and function yields. Second coroutine gets in before threshold and gets to L68, sets is_deferred and sleeps on L70. First coroutine is finished and unsets on L76. Third coroutine enters still before threshold so L62 doesn’t influence it as planned and gets to L70, sleeps similar time as second coroutine. Then both, second and third coroutines call function in very short time difference.
LLM: **Potential Time of Check to Time of Use (TOCTOU) race condition in ‘Throttler.call()‘** (Lines 58-77):
- The is_deferred flag is used to determine if a function should be deferred. However, multiple threads can check the is_deferred flag almost simultaneously and hence more than one thread might execute the contained code.
- **Solution**: A synchronization mechanism (like a lock) should be employed to ensure thread safety.
Explanation: LLM indicated the security defect identified by the reviewer, with the correct type, accurate code line range and detailed description, thus categorized as Instrumental.
Helpful (H) The type or location of the identified security defect differ to some extent from the target security defect, or the description may be too broad or partially inaccurate. However, the response still provides useful evidence pointing toward the existence of the target defect. Reviewer: (Line 141) If connect_info ever includes a secret of a type other than token (e.g. password) that mask_password looks for, you’d be replacing that with the masked token instead of masking the password or whatever it is. But I guess that’s unlikely.
LLM:**Potential Information Disclosure**(Line 141):
- Logging connection information can potentially expose sensitive details. Even though ‘mask_password‘ is being used, it’s essential to validate that no other sensitive information is leaked.
- **Solution**: Limit the details being logged and ensure that all sensitive data is properly obfuscated before logging.
Explanation: The reviewer indicated that the mask_password function might incorrectly replace different types of sensitive information. While the LLM correctly identified the relevant, it only provided general emphasis on checks for potential sensitive information exposure, failing to recognize this specific risk scenario. Consequently, this instance was categorized as Helpful rather than Instrumental.
Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task. Reviewer: (Line 329) This does implicit format detection on the imported disk, which is a well-known security hole. In this specific case a malicious user could create a ’raw’ image which actually contained a malicious qcow2 header specifying a backing file of the the compute host’s root disk. This would result in copying the compute host’s root disk to the specified rbd volume and exposing it to the user.
LLM: Based on the provided code snippet, no security defects are detected. For a more comprehensive analysis, a full code audit in the context of the entire application and its environment would be recommended.
Explanation: LLM did not detect any security defects, but the reviewer actually identified a security hole in the disk format detection. Therefore, this instance was considered Misleading.
Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed. Reviewer: (Line 469) So not even timeout cancels the future? Will this cause that threads will hang around forever leaking OS resources? This feels dangerous.
LLM: **Potential Insecure Direct Object References (Lines 89-563):**
- The ‘RequestContext‘ class and its methods provide direct access to project and user identifiers, and various methods in this file allow direct manipulation of database and message queue connections. If access controls are not properly enforced or if input validation is weak, this could lead to unauthorized access or manipulation of data.
Explanation: The reviewer identified that the thread handling may lead to OS resource leaks. However, LLM did not mention any about resource leak, instead highlighting a potential improper access issue. Since access mechanisms might be uniformly managed by a middleware, and considering the unavailability of detailed implementation information, we cannot determine whether the issue pointed out by LLM exists. Thus this instance was categorized as Uncertain.
2.3.3 Evaluation Metrics
During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements. We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al. (2020) for measuring responses in Stack Overflow. We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al. (2020) because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. Specifically, each response was categorized into one of the four rating categories (Instrumental, Helpful, Misleading, Uncertain) detailed in Table 2. To ensure clearer differentiation, for the ‘Helpful’ and ‘Instrumental’ categories, we specified their definitions based on their performance across three key attributes: defect type, location, and description. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers. At the same time, a lower M-Score signifies fewer misleading results produced during security defect detection. Given that LLMs exhibit considerable non-determinism, we repeated the experiment for each LLM-prompt combination three times and calculated the average of the three performance scores as the final result, so as to alleviate this threat and ensure a reliable evaluation of LLMs’ capabilities in security defect detection.
2.3.4 Baselines
We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Marjamäki et al., 2024) for C/C++ and Bandit (PyCQA, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (Dennison et al., 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
2.3.5 Data Labelling
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. Since the responses of LLMs were provided in the form of natural language text, to rate the responses accurately, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ datasets. We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 3.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
2.3.6 Consistency Calculation
To measure the non-determinism of LLMs, we adopted the metrics in Chang et al. (2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, an LLM generated three responses for a given code file under a specific prompt template. We then calculated the entropy for the four response categories in the three responses. We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
2.4 Research Procedure of RQ2
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses. We employed the open coding and constant comparison (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
Based on the problem categories formulated by Kabir et al. (2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. In particular, we excluded the Consistency theme defined in Kabir et al. (2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, ten responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, two authors completed the remaining data analysis task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the two authors completed the remaining data analysis and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
We consider the finalized codebook to cover all types of quality problems present in the LLM-generated responses within our dataset. To further analyze the proportion of hallucinations in LLMs’ responses, we adopted the granular taxonomy proposed by Huang et al. (2025), which categorizes hallucinations into two main types: Factual and Faithfulness. Factual Hallucination refers to outputs that contradict real-world facts or lack factual grounding, including Factual Contradiction and Factual Fabrication. Faithfulness hallucination, on the other hand, refers to deviations of the output from the instruction, current context, and reasoning process, including Instruction Inconsistency, Context Inconsistency, and Logical Inconsistency. We mapped the quality problems defined in our codebook to these hallucination subtypes to calculate their distribution across the dataset. Detailed mapping and results are provided in Section  3.2.
2.5 Research Procedure of RQ3
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the non-determinism of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is described in the following subsections.
Table 3: Factors that may influence the performance of LLM in security code review
Factor Source* Definition Rationale
Token Adjusted from  Paul et al. (2021); McIntosh et al. (2016) Number of tokens of the code file. Computed using the TikToken OpenAI (2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
FileType Created Type of the code file, i.e., Source or Auxiliary. Source indicates files that directly contain the logic of the application (i.e., .cpp, .c and .py) while Auxiliary files are utilized for declarations (i.e., .h and .hpp). The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
SecurityDefectType Transformed from  Chen et al. (2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. Yu et al. (2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities Chen et al. (2023); Zhou et al. (2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit Collected from  Paul et al. (2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity Collected from  Paul et al. (2021); McIntosh et al. (2016) McCabe’s Cyclomatic Complexity McCabe (1976). Computed using the Lizard Terry (2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases Paul et al. (2021).
Community Created Name of the community source of the code file, i.e., OpenStack or Qt. OpenStack and Qt differ significantly in terms of their structure, technology stack, project function, and main components. The LLMs’ performance may differ on data derived from these two communities.
AuthorExperience Collected from  Paul et al. (2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
AnnotationRatio Created The ratio of manual annotations in the code file. If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
AnnotationHasCode Transformed from  Chen et al. (2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs Chen et al. (2023), leading to incorrect security defect detection.
AnnotationisSecurityRelated Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects). Keywords in the description related to target security defect in annotations can act as cues for LLMs, while the descriptions of non-target security defects often involve the historical fixing or security mechanisms, which may cause interference to LLMs.
    The sources of factors correspond to the descriptions in bold in Section 2.5.1.
2.5.1 Attributes Construction
To comprehensively collect factors influencing the detection of security defects by LLMs, excluding the prompt and the model itself, we undertook the following steps:
First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work. Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntosh et al. (2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens. Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al. (2023), transforming them into factors for our experiment. Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations. In the end, we obtained a list of 10 factors in Table 3, along with the source, definition, and rationale for each factor.
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. We aggregated the dataset from three repetitive experiments after filtering out code files with multiple security defects. We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment, and their factor values were ready for model fitting.
2.5.2 Regression Model Construction
Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 3. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 3, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package [^2] https://cran.r-project.org/web/packages/rms/index.html, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Figure 3: The Phi_K correlation coefficients of explanatory variables
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating Spearman’s rank correlations (${\rho}^{2}$) between them. As  Harrell (2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned higher degrees of freedom.  Paul et al. (2021) suggested limiting the maximum allocated degrees of freedom for an explanatory variable to five or fewer, thus reducing the risk of overfitting. According to the computation results shown in Fig. 4, for interval variables, we considered using the rcs function from the R’s rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
Figure 4: Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015). We tested models with and without allocating 3 degrees of freedom to the ‘Token’ factor to balance model complexity and fit. The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2025).
3 Results
3.1 Performance of LLMs (RQ1)
Based on the evaluation metrics defined in Section 2.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts. We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively. If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperformed the others. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 4 and Table 5), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 6). The best values for each score are highlighted in bold and marked with asterisks in the tables.
From Table 4 and Table 5, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 6 demonstrated that in the selected 57 cases, Llama 2 70B scored lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in the complete datasets. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools. This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions. As such, features can only be accurately modelled using dynamic analysis, and these security defects are challenging for static analysis tools to detect (Charoenwet et al., 2024).
Table 4: Performance scores and entropy for each LLM-Prompt combination on the Python dataset
Python
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 Resp. Count 123 123 123 123 122
I-Score 1.08% 1.35% 0.27% 0.27% 1.37%
IH-Score 2.71% 3.80% 1.35% 0.27% 5.19%
M-Score 89.37% 84.28% 92.41% 97.29% 69.12%
Entropy 0.2572 0.3692 0.2070 0.0746 0.6842
GPT-4 Resp. Count 175 175 175 174 174
I-Score 4.76% 3.43% 3.24% 5.17*3.06%
IH-Score 9.34% 9.52% 8.38% 14.56*8.62%
M-Score 61.31% 60.95% 46.47% 44.06% 61.69%
Entropy 0.7426 0.7283 0.8600 0.9416 0.7450
GPT-4 Turbo Resp. Count 201 201 201 201 201
I-Score 0.66% 0.33% 0.66 0.83 0.17
IH-Score 0.66 0.83 2.32 1.99 0.33
M-Score 97.18% 96.35% 85.74% 92.87% 98.01%
Entropy 0.1757 0.0746 0.3370 0.6125 0.0404*
Gemini Pro Resp. Count 162 161 160 160 159
I-Score 0.00% 0.00% 0.63% 1.25% 0.21%
IH-Score 0.41% 0.00% 0.84% 3.54% 0.21%
M-Score 90.54% 97.09% 84.04% 55.21% 98.53%
Entropy 0.0777 0.0901 0.2424 0.1873 0.0548
Llama 2 7B Resp. Count 62 62 61 61 50
I-Score 0.00% 0.54% 1.62% 2.19% 1.33
IH-Score 1.61% 1.61% 5.40% 8.74% 4.67%
M-Score 87.10% 83.33% 56.78% 40.98% 33.33*%
Entropy 0.2625 0.3070 0.4766 0.4351 0.5521
Llama 2 70B Resp. Count 62 62 62 61 47
I-Score 0.54% 1.07% 2.15% 3.83% 1.42%
IH-Score 4.84% 4.30% 6.45% 11.48% 3.55%
M-Score 79.57% 84.41% 58.60% 43.17% 36.17%
Entropy 0.4470 0.3474 0.3514 0.3832 0.5762
Baselines SAST SonarQube CodeQL Bandit Semgrep
Resp. Count 258 258 258 258
I-Score 0.00% 0.39% 0.00% 0.00%
IH-Score 0.00% 1.16% 1.16% 0.00%
M-Score 91.47% 82.56% 78.29% 90.31%
Table 5: Performance scores and entropy for each LLM-Prompt combination on the C/C++ dataset
C/C++
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 Resp. Count 141 141 141 139 137
I-Score 1.89% 2.13% 0.00% 0.24% 0.73%
IH-Score 5.67% 3.31% 0.47% 0.96% 1.46%
M-Score 76.60% 86.05% 99.29% 95.44% 84.43%
Entropy 0.4860 0.2439 0.0195* 0.1153 0.2931
GPT-4 Resp. Count 247 247 247 247 247
I-Score 4.45% 3.91% 3.91% 5.40% 5.26%
IH-Score 7.96% 8.01% 9.31% 11.74% 11.47%
M-Score 55.47% 64.24% 44.53% 39.00% 43.59%
Entropy 0.6215 0.5329 0.6505 0.6955 0.7330
GPT-4 Turbo Resp. Count 269 269 269 269 269
I-Score 0.25% 0.12% 0.25% 0.50% 0.25%
IH-Score 0.25% 0.37% 1.24% 1.61% 1.12%
M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
Gemini Pro Resp. Count 223 221 221 221 220
I-Score 0.00% 0.15% 0.30% 1.66% 0.00%
IH-Score 0.45% 0.15% 0.90% 3.77% 0.45%
M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
Llama 2 7B Resp. Count 57 57 57 51 33
I-Score 0.58% 0.00% 0.00% 2.61% 1.01%
IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
Llama 2 70B Resp. Count 57 57 57 51 37
I-Score 1.17% 1.17% 1.75% 6.54*1.80%
IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
Baselines SAST CppCheck Semgrep
Resp. Count 276 276
I-Score 0.00% 0.00%
IH-Score 0.72% 0.36%
M-Score 98.19% 96.74%
Table 6: Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 I-Score 3.51% 3.51% 0.58% 0.58% 2.34%
IH-Score 7.60% 8.77% 1.17% 0.58% 4.09%
M-Score 81.87% 79.53% 94.74% 95.32% 77.19%
GPT-4 I-Score 8.77% 8.77% 5.85% 12.87*9.94
IH-Score 20.46% 19.30% 14.04% 23.98*19.88
M-Score 43.86% 42.11% 45.61% 40.35% 41.52%
GPT-4 Turbo I-Score 0.58% 1.75% 2.34% 2.34% 0.00%
IH-Score 0.58% 3.51% 5.26% 4.09% 1.75%
M-Score 97.08% 91.23% 75.44% 88.30% 92.98%
Gemini Pro I-Score 0.58% 0.00% 0.00% 0.58% 0.00%
IH-Score 1.17% 0.00% 0.58% 2.92% 0.00%
M-Score 96.49% 100.00% 88.30% 76.61% 100.00%
Llama 2 7B I-Score 0.58% 1.17% 1.17% 0.58% 1.75%
IH-Score 1.75% 2.34% 4.09% 2.34% 7.60%
M-Score 87.13% 84.21% 60.82% 53.80% 25.73%
Llama 2 70B I-Score 2.34% 1.17% 2.92% 5.26% 4.09%
IH-Score 4.68% 2.34% 7.02% 17.54% 4.68%
M-Score 70.18% 80.12% 42.69% 20.47*44.44%
3.1.1 RQ1.1
When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4 and Table 5), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%). Following closely behind is Llama 2 70B and GPT-3.5. GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit. Similar results are observed in Table 6. Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, GPT-4 achieved I-Score = 8.77%, IH-Score = 20.46%, and M-Score = 43.86%. It is evident that GPT-4’s performance on the selected 57 instances exceeds its performance across the entire dataset. Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
3.1.2 RQ1.2
As shown in Table 4 and Table 5, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small. Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively. Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement. A possible reason is that the name of components or packages utilized may already imply the source project of the code file, thereby diminishing the impact of project information provided in $\textbf{P}_{r}$. Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific. We speculate that some LLMs may already understand the task described in the prompt quite well, and thus, introducing a persona could add unnecessary complexity to the prompt and lead to distractions.
We can see from Table 4 and Table 5 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflects that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects. Still, it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, leveraging CWE information in the prompt is worthy of consideration.
3.1.3 RQ1.3
Table 4 and Table 5 illustrate that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B in both Python and C/C++ datasets. However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs. As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets. It can be seen that the impact of supplementing missing context by CoT prompting on the performance of LLMs is multifaceted and unstable. A possible explanation is that different LLMs vary in their ability to process and understand input. Moreover, we instructed the LLM to generate missing code context based on the corresponding commit message in $\textbf{P}_{cot}$. On the one hand, the generated context may contain the critical information needed to analyze the target security defect, which can help the LLM in security code review. On the other hand, the consistency between this generated context and the actual code context cannot be guaranteed. We observed that in some cases, the code context generated under Prompt $\textbf{P}_{cot-1}$ was inconsistent with the actual code context, resulting in the original security defect no longer being identified as a defect within the generated context. Consequently, when Prompt $\textbf{P}_{cot-2}$ was applied, the detection of security defects in the provided code was misled by the generated code context and yielded erroneous results.
3.1.4 RQ1.4
According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater non-determinism.
3.2 Quality Problems in Responses (RQ2)
3.2.1 RQ2.1
Figure 5: Average proportion of each problem type present in responses generated by GPT-4
The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig. 5, where the same color represents types within the same theme, with darker shades indicating the themes. From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
Correctness: The responses reveal two types of incorrectness: Incorrect fact (6.25%) and Incorrect concept (10.13%). The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario. Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type. As shown in the example below (Example 1), the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
Example 1 - Incorrect concept
GPT-4: ...**CWE-682 (Incorrect Calculation):**
**Issue:** The method _inventory_has_changed uses a comparison logic that might not be entirely accurate in all scenarios...
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing an overly general fix suggestion. An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
Example 2 - Vague statement
GPT-4: ...
**CWE-703(Improper Check or Handling of Exceptional Conditions)**
**General Concern**: Throughout the code, there are instances where error conditions might not be handled comprehensively...
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appears in the responses generated by GPT-4. Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes. As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence it is deemed to be Unnecessary information.
Example 3 - Unnecessary
GPT-4: ...**CWE-284 (Improper Access Control)**:
- There doesn’t seem to be any direct issue related to improper access control in the given code...
Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform a security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt. The example below (Example 4) shows that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
Example 4 - Incomplete answer
GPT-4: ### 3. CWE-664
**Potential Issue:** The creation of temporary files and directories (e.g., in function createQmlrcFile‘) ... can potentially lead to resource management issues.
**Solution:** Implement thorough checks and error handling...
While Kabir et al. (2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours. Specifically, Kabir et al. (2024) requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues. In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%. This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific application areas.
3.2.2 RQ2.2
We selected all items related to hallucination from the quality problem types in RQ1.1 and mapped them to the predefined set of hallucination subtypes. The details of mapping, corresponding explanations, and proportion results are shown in Table 7. Since a single response may exhibit multiple types of hallucination simultaneously, the sum of the proportions for each subtype exceeds the proportion of responses containing at least one hallucination. Overall, 20.19 % of the responses contain hallucinations. Among these, Factual Contradiction is the most frequent subtype, accounting for 10.13%, followed by Context Inconsistency and Factual Fabrication, which appear in 6.25% and 5.62% of responses, respectively. To alleviate these hallucinations, Retrieval-Augmented Generation (RAG) can be employed to incorporate external factual information (Tonmoy et al., 2024; Lewis et al., 2020), and post-training can also be applied to enhance the model’s capability to capture and retain contextual knowledge (Kumar et al., 2025).
Table 7: Mapping of hallucination and quality problem types with corresponding proportions.
Halluc. Type* Subtype QP Type* Explanation Prop.*
Factuality Hallucination Factual Contradiction Incorrect Concept In Incorrect Concept cases, the model’s output contradicts the authoritative definition of the corresponding CWE, thus categorized as Factual Contradiction. 10.13%
Factual Fabrication Groundless Speculation Groundless speculation refers to subjective assumptions made without contextual evidence or knowledge support. As it involves fabricated factual content, it is categorized as Factual Fabrication. 5.62%
Faithfulness Hallucination Instruction Inconsistency Detection Failure Detection Failure refers to cases where the model fails to perform the security checks or misinterprets the task, leading to deviations from the instruction. Therefore, it is categorized as Instruction Inconsistency. 0.15%
Context Inconsistency Incorrect Fact Incorrect Fact denotes cases where the line numbers, identifiers, or code content generated by the model do not align with the provided code context, thus categorized as Context Inconsistency. 6.25%
Logical Inconsistency Logical Contradiction Logical Contradiction refers to cases where the model’s reasoning about a security defect is internally contradictory, thus classified as Logical Inconsistency. 1.11%
Any Hallucination 20.19%
    Abbreviations: Halluc. = Hallucination, QP = Quality Problem, Prop. = Proportion.
3.3 Factors Influencing LLMs (RQ3)
3.3.1 Model Stability
To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the non-determinism of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
Table 8: Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(>Chi)
Token 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
Community Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
FileType Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
AnnotationisSecurityRelated -1:0 0.97 2 9.95 0.0069**
1:0 2.69
AuthorExperience 0.87 1 4.89 0.0271*
Commit 1.01 1 3.06 0.0801
SecurityDefectType Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
Crash:Race Condition 0.86
Deadlock:Race Condition 0.92
DoS:Race Condition 0.87
Encryption:Race Condition 1.18
Improper Access:Race Condition 0.68
Integer Overflow:Race Condition 0.86
Resource Leak:Race Condition 1.41
Use After Free:Race Condition 0.37
Extra:Race Condition 1.43
Complexity 1.03 1 1.74 0.1868
AnnotationRatio 0.56 1 1.55 0.2126
AnnotationHasCode 1.11 1 1.09 0.2973
  • Statistical significance of explanatory power according to Wald $\chi^{2}$ test:
    ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$
3.3.2 Explanatory Power of Each Factor
The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable. A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model. We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact. As shown in Table 8, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (<.0001). Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$<.0001) and ($\chi^{2}$=18.28, $p$<.0001). Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively. However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold. Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom. As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
3.3.3 Relationships between Factors and Responses
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 8). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files with fewer tokens, that contain functional logic, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations, the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that annotations may disrupt the LLM and are less likely to detect the target security defect.
4 Discussion
A multi-layered review strategy is suggested to apply LLMs to security code review. As the results of RQ1 indicate, the existing general-purpose LLMs still fall significantly short of reaching the effectiveness of manual security code review. However, LLMs can serve as auxiliary tools to assist reviewers. The findings of RQ3 suggest that LLMs are more adept at identifying relatively simple security defects in shorter, logic-centric code written by developers with less project experience. We recommend a multi-layered review strategy to fully leverage the capability of LLMs - reviewers can first use LLMs to conduct an initial automatic review of those short and simple commits for security defects, then review both the submitted code and the LLM’s detection results to improve efficiency. This would allow reviewers to concentrate on an in-depth review of those complex commits, including lengthy code written by more experienced developers, thereby reducing the risk of missing sophisticated security defects. Additionally, where LLMs currently show limitations, these complex commits should be more heavily incorporated into datasets used for fine-tuning LLMs to improve their effectiveness in security code review.
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy. However, in real-world security reviews, although with experience and domain knowledge, the reviewers lack prior knowledge of all potential types of security defects that may be present in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al. (2023), providing a list of all 75 CWEs in their prompts decreased the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects. Due to the limited context length of the used LLMs, inputting the entire codebase during prompt design was not feasible. The lack of code context made it difficult for the LLM to fully understand the dependencies and interaction logic between files, which in turn affected its performance in identifying security defects, such as Race Condition and Insecure Library Version. We tested various approaches, including one-shot and few-shot prompting strategies, as well as providing the complete patchset within the prompt. These strategies required splitting the code into sections for input, which we found significantly degraded the performance of LLMs compared to the basic prompt $\textbf{P}_{b}$. In order to provide as much effective information as possible within the limited context window, we designed prompts that combined multiple elements, including project metadata, CWEs, code diffs, and the commit messages. We also standardized the output format to help the model better understand the task and introduced CoT reasoning to guide the model in inferring missing context. Ultimately, we identified and selected four prompts that outperformed others, which, together with the basic prompt, were selected as templates for our formal experiments. The results demonstrate that including a list of CWEs significantly enhances LLM performance, making it our study’s most effective prompting strategy. As the context window of LLMs continues to expand and prompt engineering techniques evolve, we plan to explore additional strategies in the future, such as incorporating RAG for external information and employing automated prompt optimization to dynamically adjust prompts during reasoning, thereby further improving the capability of LLMs in security defect detection.
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better supporting LLMs in code review tasks.
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amounts of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo having a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2. Therefore, ensuring the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results. The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses. This may be because the outputs generated by LLMs with higher non-determinism are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower non-determinism. Both detection capability and the stability of detection results are crucial for applying LLMs in security code review. However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results. Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the non-determinism inherent in LLMs represents a valuable direction for future improvements.
5 Threats to Validity
Internal Validity refers to whether an observed association between two variables can be attributed to a causal link between them. One of the internal validity threats could be data leakage. The data used in our study was obtained from two large-scale open-source projects, which may overlap with the data used for training LLMs, a common problem with LLM-based studies. Since the training data cutoff dates for the models used (GPT-3.5 and GPT-4) are not disclosed and the dataset we used  (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time. Moreover, according to Cao et al. (2024), contaminated data does not always affect the results, and the model sometimes even performs better after the model’s cutoff date. Another threat is the design of prompts. We design our prompt according to best practices to evoke the best model performance. We selected words that are as unbiased, non-suggestive, and unambiguous as possible. We iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information. To specifically analyze its impact, we designed a prompt $\textbf{P}_{cot}$, in which LLMs are instructed to generate code context according to the commit message before detecting security defects.
External Validity concerns whether a causal link generalizes across contexts. In this study, the experiments were conducted on 6 LLMs based on 4 open-source projects across programming languages including Python and C/C++. We acknowledge that the dataset we utilized may not fully represent industrial projects across various programming languages, and the performance of the selected LLMs may not represent the capabilities of all existing LLMs. We implemented the following strategies to ensure the generalizability of the experimental results. For model diversity, our experiments incorporated both open-source and proprietary LLMs of varying parameter scales to encompass a diverse range of popular LLMs. For dataset construction, we collected code review data from real-world open-source projects in the OpenStack and Qt communities. These projects differ in architectures, programming languages, and community structures, providing a comprehensive and valuable dataset for empirical analysis.
Construct Validity assesses whether a specific set of metrics corresponds to what they are intended to measure. To measure the performance of LLMs, we classified the natural language responses of LLMs into four predefined categories, with human reviews of security defects in code as the ground truth. There is a potential natural language understanding bias which affects classification accuracy. To reduce the bias, we established granular definitions for each category in three dimensions: defect type, location, and description, supplemented by examples for each category to ensure that the metrics are accurate, appropriate, and consistent throughout the evaluation process. In addition, the data labeling in RQ1 and the analysis of quality problems in RQ2 were all carried out manually. In order to further reduce the impact of subjective judgments, we employed open coding and constant comparison (Glaser, 1965). In the data labeling employed for RQ1, the first and third authors conducted a pilot data labeling and completed all the labeling work after reaching a consensus on the criteria through discussion. In the data analysis of RQ2, the first and seventh authors, conducted two rounds of pilot data analysis to determine the definition of each quality problem and achieved a high level of consistency in the results. All discrepancies in the manual work of our study were reviewed and resolved through negotiation and discussion among at least three authors, thus ensuring the reliability of the findings.
Conclusion Validity refers to whether the conclusions drawn about relationships in the data are statistically sound and reasonable. Due to the non-determinism of LLMs, the same input may yield different outputs across different runs. Therefore, a single round of experiments is insufficient to accurately reflect the performance of LLMs. To enhance the reliability and representativeness of our evaluation, we repeated the RQ1 experiment three times. For RQ2, we calculated the average frequency of each problem across these three experiments. For RQ3, we combined the responses from the three experiments to fit the model. We also assessed their non-determinism as part of the overall performance. Specifically, in RQ1.4, we calculated the consistency of LLMs’ responses across three runs to compare the non-determinism of different LLMs under different prompts.
6 Related Work
6.1 Security Defect Detection in Code Review
Several studies have analyzed security defect detection during code reviews. Paul et al. (2021) built a dataset from the Chromium OS project, consisting of 516 code reviews that successfully identified security defects and 374 code reviews that missed security defects. The study analyzed 18 code review attributes, including review duration, reviewer experience, and author experience, to explore factors influencing the identification of security defects. It was found that the duration of code reviews and the frequency of peer reviews positively affected the detection of security defects, while the number of directories under review and the number of prior commits had a negative impact. Alfadel et al. (2023) examined pull requests (PR) from 10 popular JavaScript GitHub projects within the npm ecosystem, then analyzed the frequency, categories, and handling of security defects identified during code reviews. Their work indicated that only a small fraction of PRs raise security problems, and most of the identified security problems were resolved during the code review.
6.2 Code Characteristics of Security Defects
Security defects mainly stem from improper system design or low-quality code, caused by a lack of security expertise or violations of secure coding practices during implementation. Many studies have focused on analyzing code for vulnerabilities and security defects in hopes of helping developers understand how security problems originate in a system, so that defects are detected early.
Alves et al. (2016) constructed a security defect dataset based on five widely used projects and 27 software metrics, such as lines of code, cyclomatic complexity and knots in the control flow graph, among others. This study demonstrated that most of these metrics are effective in distinguishing vulnerable from non-vulnerable code, thereby providing valuable support for security practitioners. Meneely et al. (2013) conducted quantitative and qualitative analyses on 124 Vulnerability-Contributing Commits (VCC) in the Apache HTTP Server to identify characteristics of code changes with security defects. The findings suggest that longer lines of code and less experienced committers are more likely to introduce security defects. Bosu et al. (2014) identified 413 VCCs from code reviews in 10 open-source projects and reported consistent observations with Meneely et al. (2013) that longer lines of code and less experienced committers are more likely to introduce security defects. Additionally, they indicated that modified code files are more likely to contain security defects than new files.
6.3 Automated Tools of Security Defect Detection
Various tools have been developed to automatically detect security defects. Since our study focuses on using LLMs to assist reviewers in identifying security defects during code review, we limit our coverage to studies related to security defect detection, excluding those studies on security defect remediation or program repair.
Charoenwet et al. (2024) conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of VCCs. By combining several tools, warnings could be produced for 78% of the VCCs. However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review. Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
Several studies have explored the performance of LLMs in detecting security defects. The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type. Zhou et al. (2024b) utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgment under various prompts. They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art detection approaches, while GPT-4 consistently outperformed the state-of-the-art. Purba et al. (2023) compared the performance of LLMs with popular static analysis tools in binary detection. Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools. Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description, and suggestions for defect remediation.
Previous studies also investigated whether LLMs could provide more information in security defect detection. Bakhshandeh et al. (2023) provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers. Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools. Yin and Ni (2024) evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description, and found that the existing state-of-the-art approaches, such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023), generally outperformed LLMs. Specifically, LLMs exhibited limited overall accuracy and varied performance across different CWE types in defect localization and description tasks. These studies have all been limited to quantitatively evaluating LLM performance. In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to understand the potential and challenges LLMs face in security code review.
Employing specific prompting frameworks to enhance the performance of LLMs in identifying security defects has also been explored. Yang et al. (2024) introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection. Their framework yielded superior performance compared to state-of-the-art prompting techniques. Nong et al. (2024a) proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution). The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours. These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements. Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
It is worth noting that the studies mentioned above focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs. Inspired by these related works, we aim to leverage LLMs’ potential in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.
7 Conclusions and Future Work
In our study, we comprehensively analyzed the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems in the responses generated by this LLM. We performed a regression analysis to explore factors influencing its performance.
Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to state-of-the-art static analysis tools. However, their application remains limited. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) During code review, GPT-4 is more effective in identifying security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project. (4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.
We plan to extend this work in the next step by incorporating code reviews from proprietary software projects into our dataset and attempting enhancement strategies such as few-shot learning and fine-tuning. In addition to applying LLMs to help reviewers detect security defects during code review, we also plan to analyze the ability of LLMs to fix security defects in code so that developers can comprehensively evaluate the potential of LLMs in the task of secure code review.
Data availability
We have shared the link to our dataset in the reference (Yu et al., 2025).
Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.
[…]